Theta Analytics is currently traded at $27.28 per share. Theta anticipates paying a cash dividend of $2.71 at the end of the year, with that amount growing at 3.6 percent per year forever after. What is the cost of equity capital in percent?
Answer #1

Current stock price (P0) = $27.28

Dividend after year 1 (D1) = $2.71

Dividend growth rate (g) = 3.6%

Cost of equity (Ke) = ?

Ke = D1/P0 + g

= 2.71/27.28 + 0.036

= 0.0993 + 0.036

= 0.1353

= 13.53%

Cost of equity (Ke) = 13.53%

Kindly give a positive rating if you are satisfied with the answer. Feel free to ask if you have any doubt. Thanks.